OpenStudy (anonymous):

factor the expression given below write your answer in the form (r+a)(s+b) where r,s,a and b can be any numbers or expressions 4x^2-y^2z^2

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

Okay if you look at this we have 2x*2x-yz*yz, we have to figure out how to write this in factored form. we already know the first term is going to be 2x, because 2x*2x=4x^2 (2x )(2x ) now we already figured out that y^2z^2 is the same as yz*yz, so the second term will be yz so we hve (2x yz)(2x yz) but since the answer is 4x^2-y^2z^2 there must be one + sign and one - sign so that it doesnt become positive so it is (2x+yz)(2x-yz)

OpenStudy (anonymous):

ok i have 5 more questions can u help

OpenStudy (anonymous):

I have to make lunch, sorry

OpenStudy (anonymous):

the polynomial x^2+16x+100 can be written as (x+a)^2+b where A.a=8 and b=36 B, a=-8 and b-18 C,a=6 and b=18 D. a=36 and b=8

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

can someone help with the question above

OpenStudy (anonymous):

@ganeshie8